# Image as an IMU: Estimating Camera Motion from a Single Motion-Blurred Image

### v1

@@ 1 Introduction
A new approach for estimating camera motion that uses the blur present in a frame to provide extremeley robust and accurate relative pose estimates
@@ 2 Related Works
Motion from Blur. This paper is most closely related to the line of work referred to as motion from blur, which aims to extract pixel flow information only from a single motion blurred image. [5] [26] are among the earliest works within this topic, where [5] introduced a linear model for computing the motion vectors and [26] independently analyzed the frequency domain to identify blurred regions. However, both methods assume that there is only linear, spatially-invariant blur in the image. [32] built on top of [26] by using the three color channels of an image. The seminal work [7] proposed a straightforward method where flow could be estimated using the alpha channel of an image. A limitation of these aforementioned methods is the inability to distinguish directional information with the motion vectors. [11] resolves this and introduces a variational approach to estimate the motion vectors and the blur segmentation in isolation. [1] proposed the first deep learned solution using a spatial transformer network. These methods only focus on estimating a dense flow field from the motion blurred image, whereas we are interested in the usefulness of the flow field for obtaining camera motion.
@@
Our method also estimates the depth of the scene given the motion blurred image to retrieve the metric camera motion. Similarly [16], [20], and [25] demonstrate how to recover the 3D geometry even in motion-blurred scenes. [20] is the most relevant to our work and presented a visual odometry method which computes the homography during exposure time of a blurred image, which demonstrates increased robustness in severely blurred videos. However, they minimize the photometric error between a sharp keyframe and a blurred frame, where they use a off-the-shelf deblurring method to do so. We do not require any deblurring and only need a single image to estimate the motion.
@@
Camera Pose Estimation. Camera pose estimation is a fundamental computer vision task used in 3D reconstruction which can be categorized into SfM and SLAM/VO methods. SfM methods operate on an unordered set of images and often involve estimate the camera intrinsics and extrinsincs. SLAM/VO methods can be viewed as a subset of SfM with the added constraint for running in real-time, with and without loop closures respectively.
@@
For SfM methods, COLMAP [31] [30] is an incremental-SfM method that is commonly the default choice for estimating camera parameters, poses, and the scene geometry. ORB-SLAM [3] and DSO [9], on the other hand, are classic real-time methods for obtaining camera poses and mapping the environment. While highly accurate in sequences with large overlap, these traditional systems succumb to failures under adversarial conditions, including few-view reconstruction and motion blur. These failures motivate deep-learning based approaches which are more robust in these difficult scenarios. Several fully-differentiable SfM pipelines have been proposed for greatly improving 3D reconstruction and pose estimation robustness [42] [34] [2]. In particular, DUSt3R [44] and MASt3R [17] have shown that directly regressing scene geometry in the form of pointmaps enables incredible performance, which has inspired a line of “-3R”-related works [48] [41] [33] [46] [43].
Similarly, SLAM and VO methods have also underwent a deep-learning revolution and have similarly adopted learned modules with improved performance [6] [45] [37] [35] [28] [24]. For example, DROID-SLAM [37] is an end-to-end solution for obtaining accurate real-time reconstructions using RAFT [36] for estimating optical flow. These approaches rely on using two or more images for estimating the camera motion, whereas our method is able to use a single motion-blurred frame to estimate camera velocity. Due to there being no existing methods that estimate camera motion from a single frame, we opt to use these standard camera pose estimation methods, specifically MASt3R, DROID-SLAM and COLMAP, and estimate the velocities with a finite-difference approximation.
@@ 5.1 Synthetic Dataset
To train our model, we require a dataset of blurred images with corresponding ground-truth flows, depths, and relative poses across exposure time. Since no public dataset meets these requirements, we create our own by adapting a subset of ScanNet++v2 [47]. Our data pipeline is shown in Figure 3.
@@
We only use the first and last virtual frames and for obtaining the ground-truth flow field. We use the low-resolution ARKit depth and in PromptDA [18] to obtain our dense, high-resolution depth map as ground truth. For all pixels, we then perform:
which uses to backproject all pixel points in and subsequently project them into . Afterwards, we retrieve the flow field with the displacements of the projected points as expressed in Equation 3.
@@ 5.2 Real-world Dataset
As our method, including the velocity solver, is fully-differentiable we can train it in an end-to-end manner. This enables us to additionally train on real-world motion blur to further close the reality-gap between our synthetically blurred dataset and in-the-wild images. We therefore collect 10k real-world motion-blurred images with corresponding ARKit poses, IMU measurements, and exposure times across about 30 scenes. We follow Liu et al. [21] and perform the Fast Fourier Transform to identify blurred images, where images below a specified threshold are blurry enough to be used for training or validation. We also supplement with a small amount of motionless images (with no blur) to enhance the robustness to still frames.
@@ 6 Experiments
Our model is trained in three stages. We first train the flow and depth decoders without pose supervision on our synthetically blurred dataset with a batch size of 32. Pose supervision is added once the depths and flow estimates are reasonable, and train poses, flow and depths with a batch size of 8 for 300k steps. Finally, we finetune the model on the real-world collected images to close the reality gap from synthetic images for 10k steps. We train our method using the Adam optimizer [15] on a single Nvidia RTX 3090.
@@
To assess the accuracy of our method, we evaluate the rotational velocities and translational velocities in four real-world videos (see Supplementary 2). For each video sequence, we get the ARKit-estimated camera poses, IMU measurements, and the exposure times for each image. We directly compare predictions to the gyroscope measurements and predictions to a finite-difference velocity approximation from the ARKit translations. We treat the ARKit translations as the ground-truth as their visual-inertial method fuses additional sensor information, which helps mitigate the poor visual input. We measure the RMSE across each axis for and .
@@
Results.
We report the rotational and translational RMSE in Tables 1 and 2. While the recorded video sequences suffer from severe motion blur, our method is able to handle the motion very robustly. Figure 5 highlight the overall robustness of our method, with the error CDFs demonstrating that our method suffers from noticably less errors than the baselines. In these sequences, the baseline methods suffer from major outliers and zero-prediction failures, which can be observed in Figure 4. Velocity predictions from the billiards sequence are shown with the remaining sequences included in Supplementary 3. Overall, we obtain an average of 31% reduction from the second-best result in rotational velocities and a 24% reduction in translational velocities.

### v3

@@ 1 Introduction
A new approach for estimating camera motion that uses the blur present in a frame to provide extremely robust and accurate relative pose estimates
@@ 2 Related Works
Motion from Blur. This work is related to motion from blur methods that attempt to extract pixel-wise flow from a single blurred image. Early approaches assumed spatially invariant blur, with [7] using a linear motion model to extract the flow and [35] using a frequency domain approach. Later methods extended these where [10] first extracts an alpha channel matte and uses the edge gradients to estimate object motion, [41] builds on [35] to use the color channels directly while [14] proposed a variational technique and [1] a learned approach. Our focus, however, is not on estimating the flow field but rather the actual camera motion.
Several works aim to recover either rotational or translational camera motion directly from a blurred image. A seminal work in this direction is [21] which proposes to estimate the camera rotation directly from a single motion-blurred image by searching for edgels to estimate the rotation axis, which is used to help compute the blur length and subsequently the rotation. [24] [3] also estimate the camera rotation with assuming negligible translations, while [2] [8] estimate exclusively linear motion. In contrast, we estimate the full 6DoF velocity. Closer to our method, [62] proposes a neural network to estimate the full relative pose from a single motion-blurred image but directly regresses the pose parameters and requires a clean depth map as input.
@@
Our method estimates depths to retrieve the metric-scaled motion. Similarly, [22], [27], and [34] recover 3D geometry in motion-blurred scenes. [27] is the most relevant, being a visual odometry method which computes the homography during exposure time between a sharp and potentially blurry frame. However, this depend on two images and an off-the-shelf deblurring method, whereas we do not require any deblurring for a single input image.
Deblurring. An adjacent line of work is deblurring, which aims to recover a sharp image given a blurred one. [17] [33] [32] use classical energy minimization frameworks to obtain depth and camera pose estimates, similarly to our method, in addition to the deblurred image. However, compared to deep-learned deblurring methods [60], classical methods are frequently slower during test time and underperform in more general conditions. Therefore, we do not benchmark against such methods given the restrictive assumptions made and significant runtime slowdown. Several deep deblurring methods [45] [15] [61] share parallels to our work in also estimating a motion flow field. [56] reformulates removing the rolling-shutter effects as a deblurring problem with rolling shutter global reset. While our motion field estimates could in theory compute a deblur kernel, our goal is to exploit motion cues from blur artifacts rather than remove them.
@@
Camera Pose Estimation. Camera pose estimation is a fundamental task for 3D reconstruction which can be categorized into SfM and SLAM/VO methods. SfM methods operate on unordered images and estimate the camera intrinsics and extrinsics. SLAM/VO methods can be viewed as a subset of SfM with enforced real-time constraints.
@@
COLMAP [40] [39] is an incremental-SfM method that is commonly the default choice for estimating camera parameters and scene geometry, while ORB-SLAM [5] and DSO [12] are classic real-time alternatives. While highly accurate in sequences with large overlap, these traditional systems succumb to failures under conditions including few views and motion blur. These failures motivate deep-learning based approaches which are more robust in these difficulties. Several fully-differentiable SfM pipelines have been proposed for greatly improving 3D reconstruction and pose estimation accuracy and robustness [52] [43] [4]. In particular, DUSt3R [54] and MASt3R [23] have shown that directly regressing pointmaps enables incredible performance, which has inspired a line of “-3R”-related works [59] [51] [42] [57] [53]. SLAM and VO methods have also underwent its own deep revolution [9] [55] [47] [44] [37] [31], with DROID-SLAM [47] obtaining accurate SOTA pose estimates by using RAFT [46] for estimating optical flow. Since no prior methods estimate camera motion from a single image, we decide to benchmark against COLMAP, MASt3R, and DROID-SLAM, approximating velocities via finite differences.
@@ 5.1 Synthetic Dataset
To train our model, we require a dataset of blurred images with corresponding ground-truth flows, depths, and relative poses across exposure time. Since no public dataset meets these requirements, we create our own by adapting a subset of ScanNet++v2 [58]. Figure 3 shows our data pipeline.
@@
We only use the first and last virtual frames and for obtaining the ground-truth flow field. We use the low-resolution ARKit depth and in PromptDA [25] to obtain our dense, high-resolution depth map as ground truth. For all pixels, we then perform which uses to backproject all pixel points in and subsequently project them into . Afterwards, we retrieve the flow field with the displacements of the projected points as expressed in Equation 3.
@@ 5.2 Real-world Dataset
As our method, including the velocity solver, is fully-differentiable we can train it in an end-to-end manner. This enables us to additionally train on real-world motion blur to further close the reality-gap between our synthetically blurred dataset and in-the-wild images. We therefore collect 10k real-world motion-blurred images with corresponding ARKit poses, IMU measurements, and exposure times across about 30 scenes. We follow Liu et al. [28] and perform the Fast Fourier Transform to identify blurred images, where images within a threshold are blurry enough to be used. We also supplement with a small amount of images without blur to enhance the robustness to still frames.
@@ 6 Experiments
Our model is trained in three stages. We first train the flow and depth decoders without pose supervision on our synthetic dataset with a batch size of 32. Pose supervision is added once the depths and flow estimates are reasonable, and we train poses, flow and depths with a batch size of 8 for 300k steps. Finally, we finetune the model on the real-world collected images to close the reality gap from synthetic images for 10k steps. We train our method using Adam [20] on one Nvidia RTX 3090.
@@
To assess the accuracy of our method, we evaluate the rotational velocities and translational velocities in four unseen, real-world videos (see Supplementary 2). For each video sequence, we get the ARKit-estimated camera poses, IMU measurements, and the exposure times for each image. We directly compare predictions to the gyroscope measurements and predictions to a finite-difference velocity approximation from the ARKit translations. We treat the ARKit translations as the ground-truth as their visual-inertial method fuses additional sensor information, which helps mitigate the poor visual input. We measure the RMSE across each axis for and .
@@
Results.
We report the rotational and translational RMSE in Tables 1 and 2. While the recorded video sequences suffer from severe motion blur, our method is able to handle the motion very robustly. Figure 5 highlight the overall robustness of our method, with the error CDFs demonstrating that our method suffers from noticably less errors than the baselines. In these sequences, the baseline methods suffer from major outliers and zero-prediction failures, which can be observed in Figure 4. Velocity predictions from the billiards sequence are shown with the remaining sequences included in Supplementary 3. Overall, we obtain an average of 31% reduction of RMSE from the second-best result in rotational velocities and a 24% reduction in translational velocities.